Rewrite the equation as a function of x.
-3x+4y=11

Answer by benni1013(206)

Usually you will see functions solved as a function of y in the form of y= now we just need to turn it into the form of x=
-3x+4y=11
First let's move 4y to the other side by subtracting
-3x=11-4y
Last divide by -3 to get
x=-11/3+4/3y